# Balanced Mixed-Type Tabular Data Synthesis with Diffusion Models

### v1

@@ 5.4 Computational Results
This section presents the computational results of our evaluation of synthetic tabular data generated with five different random seeds. To ensure the robustness of our model-based evaluations, we employed a diverse set of classic classifiers, including decision trees, random forests, support vector machines, multilayer perceptrons, and XGBoost. We determined the optimal model for each synthetic dataset through meticulous training and evaluation across separate training and validation sets, utilizing five random seeds for each model. The averaged validation accuracy, along with the standard deviation, as shown in Table 2, revealed XGBoost as the dominant performer across most datasets, with random forest excelling in one exceptional case (see fairness scores in Section A.3).
Table 2: The best and average values of validation accuracy on experimental datasets (the averaged validation accuracy $\pm$ the standard deviation)
Dataset Model Choice Best Average
AD XGBoost $0.866_{\pm 0.000}$ $0.795_{\pm 0.041}$
BA XGBoost ${0.900_{\pm 0.000}}$ $0.869_{\pm 0.031}$
CA XGBoost ${0.733_{\pm 0.000}}$ $0.658_{\pm 0.015}$
CR Random Forest ${0.816_{\pm 0.000}}$ $0.757_{\pm 0.018}$
DE XGBoost ${0.941_{\pm 0.000}}$ $0.929_{\pm 0.001}$
KD XGBoost ${0.956_{\pm 0.000}}$ $0.946_{\pm 0.000}$
LA XGBoost ${1.000_{\pm 0.000}}$ $0.942_{\pm 0.004}$
Performance. In our exploration of the machine learning efficiency of synthetic data compared to baselines, we employ two rigorous protocols. The first protocol calculates the average machine learning efficiency across all classifiers, while the second identifies and utilizes the best classifier for each dataset. This multifaceted approach ensures a robust and insightful evaluation. As illustrated in Table 3, our proposed method maintains competitiveness in terms of average machine learning efficiency, even when incorporating a uniformly distributed direction for sensitive features during the sampling phase. It generally outperforms CTGAN and TVAE. Moreover, our method excels against TabDDPM on the majority of datasets. While the interpolation-based method, SMOTE, proves highly competitive and challenging to outperform, our model achieves a machine learning efficiency very close to SMOTE, underscoring its effectiveness. The best machine learning efficiency exhibits a similar pattern with the average machine learning efficiency, with SMOTE and our model emerging as the two most competitive methods.
Table 3: The values of machine learning efficiency averaged on all types of classifiers
Average Accuracy
AD BA CA CR DE KD LA
TabDDPM $0.777_{\pm 0.093}$ $0.874_{\pm 0.017}$ $0.640_{\pm 0.025}$ $0.768_{\pm 0.032}$ $0.920_{\pm 0.001}$ $0.937_{\pm 0.001}$ $0.905_{\pm 0.012}$
CTGAN $0.742_{\pm 0.148}$ $0.873_{\pm 0.019}$ $0.600_{\pm 0.061}$ $0.746_{\pm 0.058}$ $0.867_{\pm 0.050}$ $0.931_{\pm 0.014}$ $0.882_{\pm 0.112}$
TVAE $0.786_{\pm 0.023}$ $0.854_{\pm 0.026}$ $0.613_{\pm 0.058}$ $0.784_{\pm 0.021}$ $0.881_{\pm 0.051}$ $0.938_{\pm 0.008}$ $0.919_{\pm 0.119}$
SMOTE $0.807_{\pm 0.038}$ $0.880_{\pm 0.022}$ $0.632_{\pm 0.074}$ $0.770_{\pm 0.044}$ $0.928_{\pm 0.022}$ $0.939_{\pm 0.013}$ $0.930_{\pm 0.124}$
Ours $0.803_{\pm 0.022}$ $0.880_{\pm 0.011}$ $0.629_{\pm 0.001}$ $0.734_{\pm 0.045}$ $0.926_{\pm 0.004}$ $0.932_{\pm 0.004}$ $0.921_{\pm 0.003}$
Best Accuracy
AD BA CA CR DE KD LA
TabDDPM $0.834_{\pm 0.004}$ $0.895_{\pm 0.003}$ $0.708_{\pm 0.003}$ $0.816_{\pm 0.002}$ $0.937_{\pm 0.003}$ $0.945_{\pm 0.001}$ $0.964_{\pm 0.002}$
CTGAN $0.817_{\pm 0.002}$ ${0.888}_{\pm 0.002}$ ${0.652}_{\pm 0.006}$ ${0.716}_{\pm 0.009}$ ${0.870}_{\pm 0.008}$ ${0.941}_{\pm 0.001}$ ${0.975}_{\pm 0.002}$
TVAE ${0.810}_{\pm 0.004}$ ${0.849}_{\pm 0.008}$ ${0.662}_{\pm 0.006}$ ${0.773}_{\pm 0.006}$ ${0.900}_{\pm 0.003}$ ${0.944}_{\pm 0.001}$ ${0.993}_{\pm 0.001}$
SMOTE $0.853_{\pm 0.000}$ $0.898_{\pm 0.000}$ $0.704_{\pm 0.000}$ $0.817_{\pm 0.001}$ $0.941_{\pm 0.000}$ $0.949_{\pm 0.000}$ $1.000_{\pm 0.000}$
Ours $0.838_{\pm 0.001}$ $0.892_{\pm 0.003}$ $0.709_{\pm 0.004}$ $0.806_{\pm 0.004}$ $0.942_{\pm 0.003}$ $0.930_{\pm 0.003}$ $0.995_{\pm 0.004}$
Privacy. As explained in Section 5.3, we employ DCR values to assess the privacy score of the synthetic tabular data. The median DCR values, obtained across five random seeds, are detailed in Table 4. Remarkably, our method consistently attains a superior privacy score in comparison to both SMOTE and TabDDPM. This suggests that our approach excels in maintaining the distinctiveness and privacy of the synthetic data.
Table 4: The median value of distance to the closest record
AD BA CA CR DE KD LA
TabDDPM $0.179$ $0.195$ $0.021$ $0.123$ $0.520$ $2.089$ $0.176$
CTGAN $0.340$ $0.303$ $0.028$ $0.202$ $0.428$ $2.451$ $0.276$
TVAE $0.253$ $1.415$ $0.032$ $0.208$ $0.443$ $2.018$ $1.414$
SMOTE $0.170$ $0.191$ $0.021$ $0.117$ $0.398$ $2.005$ $0.178$
Ours $0.511$ $0.229$ $0.023$ $0.255$ $0.430$ $2.831$ $0.216$
Fairness. A key contribution of our work is the generation of balanced tabular data while considering sensitive features. To substantiate our claim and provide empirical evidence, we conduct a thorough assessment of the fairness of our proposed model. After choosing the best-performing classifier for each dataset, we assess its fairness by evaluating two metrics: demographic parity ratio and equalized odds ratio, and the results are presented in Table 5. Our approach outperforms baseline methods notably in terms of equal allocation, as evidenced by higher values of demographic parity ratio across nearly all datasets. For instance, on the KDD Census dataset, our method achieves a demographic parity ratio of $0.613$, surpassing baseline methods like SMOTE and TabDDPM by around fifty percent, while SMOTE only attains $0.141$, and TabDDPM reaches $0.097$. Similarly, on the UCI Adult dataset, our model demonstrates a demographic parity ratio of $0.529$, surpassing the ratios of $0.306$ by SMOTE and $0.312$ by TabDDPM. In addition, our model demonstrates competitiveness in terms of equal performance metrics, as indicated by the equalized odds ratio. Particularly noteworthy are datasets such as UCI Adult and KDD Census, where our model achieves an equalized odds ratio of $0.641$ and $0.884$, respectively, surpassing baseline models that register less than $0.250$ in these metrics. Furthermore, employing equitable synthetic data to train classifiers can yield higher fairness scores compared to training with the original data, as illustrated in Table 7. In contrast, baseline models generally achieve comparable fairness scores.
Table 5: The values of demographic parity ratio and equalized odds ratio with the best-performing classifier
Demographic Parity Ratio
AD BA CA CR DE KD LA
TabDDPM $0.312_{\pm 0.033}$ $0.590_{\pm 0.070}$ $0.957_{\pm 0.036}$ $0.791_{\pm 0.063}$ $0.634_{\pm 0.020}$ $0.097_{\pm 0.032}$ $0.858_{\pm 0.007}$
TabGAN ${0.188}_{\pm 0.039}$ ${0.749}_{\pm 0.060}$ ${0.760}_{\pm 0.027}$ ${0.915}_{\pm 0.051}$ ${0.706}_{\pm 0.021}$ ${0.112}_{\pm 0.037}$ ${0.844}_{\pm 0.006}$
TabVAE ${0.249}_{\pm 0.017}$ ${0.624}_{\pm 0.045}$ ${0.965}_{\pm 0.011}$ ${0.740}_{\pm 0.088}$ ${0.664}_{\pm 0.012}$ ${0.092}_{\pm 0.027}$ ${0.847}_{\pm 0.006}$
SMOTE ${0.306}_{\pm 0.000}$ ${0.625}_{\pm 0.000}$ ${0.975}_{\pm 0.000}$ ${0.787}_{\pm 0.000}$ ${0.671}_{\pm 0.000}$ ${0.141}_{\pm 0.000}$ ${0.851}_{\pm 0.000}$
Ours $0.529_{\pm 0.046}$ $0.839_{\pm 0.098}$ $0.963_{\pm 0.017}$ $0.828_{\pm 0.048}$ $0.706_{\pm 0.005}$ $0.613_{\pm 0.057}$ $0.862_{\pm 0.007}$
Equalized Odds Ratio
AD BA CA CR DE KD LA
TabDDPM $0.247_{\pm.0.033}$ $0.496_{\pm.0.077}$ $0.919_{\pm.0.068}$ $0.680_{\pm.0.081}$ $0.764_{\pm.0.110}$ $0.134_{\pm.0.050}$ $0.593_{\pm.0.484}$
CTGAN ${0.125}_{\pm 0.047}$ ${0.760}_{\pm 0.074}$ ${0.669}_{\pm 0.038}$ ${0.895}_{\pm 0.048}$ ${0.815}_{\pm 0.062}$ ${0.178}_{\pm 0.067}$ ${0.611}_{\pm 0.172}$
TVAE ${0.205}_{\pm 0.023}$ ${0.621}_{\pm 0.046}$ $0.950_{\pm 0.027}$ ${0.667}_{\pm 0.115}$ ${0.825}_{\pm 0.067}$ ${0.126}_{\pm 0.046}$ ${0.562}_{\pm 0.175}$
SMOTE ${0.231}_{\pm 0.000}$ ${0.593}_{\pm 0.000}$ ${0.939}_{\pm 0.000}$ ${0.683}_{\pm 0.000}$ ${0.827}_{\pm 0.000}$ ${0.202}_{\pm 0.000}$ ${1.000}_{\pm 0.000}$
Ours $0.641_{\pm 0.108}$ $0.760_{\pm 0.043}$ $0.901_{\pm 0.035}$ $0.802_{\pm 0.056}$ $0.254_{\pm 0.089}$ $0.884_{\pm 0.055}$ $0.170_{\pm 0.340}$
Figure 2: Comparison in SHAP values on KDD dataset: balanced scenario (top) vs. unbalanced scenario (bottom) using XGBoost
In Figure 2, we illustrate the feature importance through absolute SHAP values when making predictions with XGBoost on the KDD Census dataset. On the top, we present the scenario where synthetic data generated by our approach exhibits a uniformly distributed sensitive feature, sex, across all samples. In contrast, the figure on the bottom showcases data generated by TabDDPM with biased sampling in the sex feature. It is evident that in the data generated by TabDDPM, the variable sex significantly influences decision-making, while our method mitigates the impact of sex on the final decision to a considerable extent.
Apart from fairness evaluation dependent on machine learning models, we also analyze the class distribution within the sensitive attributes. For example, in the UCI Adult dataset, the distribution of sensitive attributes has an inherent bias. In contrast, the synthetic data generated through our approach exhibits more balanced distributions, as depicted in Figure 3, where the y-axis represents the percentage of the respective attribute. We present visualizations of class imbalance in sensitive attributes for the rest of experimental datasets in Figure 4.
@@
Figure 3: Comparison in the real vs. synthetic distribution of sensitive attributes in the UCI Adult dataset
5.5 Limitations and Discussion
Our proposed method is designed to generate balanced synthetic tabular data considering sensitive attributes. Empirically, classifiers trained on such balanced synthetic tabular data demonstrate higher fairness scores at equal allocation metrics indicated by demographic parity ratios. However, there is no guarantee that equal performance, reflected by equalized odds ratios, will always improve. We use U-Net with attention layers as the backbone neural network for the posterior estimator in the diffusion framework, but it is time-consuming. Despite our model exhibiting superior machine learning efficiency and higher fairness scores compared to other deep generative models in our experiments, the computational time is approximately five times longer. In the future, we may explore the incorporation of constraints to ensure equal performance and seek ways to reduce computational costs.

### v3

@@ 5.4 Computational Results
This section presents the computational results of our evaluation of synthetic tabular data generated with three different random seeds. We trained the SOTA machine learning model for tabular data, CatBoost, on three different sets of synthetic data, each generated with a unique random seed.
5.4.1 Fidelity and Diversity of Synthetic Data
Table 2: The values of error rates of column-wise density estimation and pair-wise column correlations.
Adult Bank COMPAS Mean Adult Bank COMPAS Mean
[…]
CoDi .$145_{\pm.000}.$$154_{\pm.000}.$$205_{\pm.001}.$ $16.8\%$$495_{\pm.000}.$$344_{\pm.001}.$$550_{\pm.001}.$ $46.3\%$
GReaT .$077_{\pm.000}.$$102_{\pm.001}.$$093_{\pm.001}.$ $9.1\%$$208_{\pm.015}.$$213_{\pm.010}.$$165_{\pm.016}.$ $19.5\%$
SMOTE .$025_{\pm.000}.$$020_{\pm.000}.$$021_{\pm.001}.$ $2.2\%$$054_{\pm.004}.$$042_{\pm.005}.$$047_{\pm.005}.$ $4.8\%$
STaSy .$102_{\pm.000}.$$182_{\pm.000}.$$108_{\pm.001}.$ $13.1\%$$163_{\pm.000}.$$221_{\pm.001}.$$138_{\pm.001}.$ $17.4\%$
TabDDPM .$037_{\pm.001}.$$028_{\pm.001}.$$057_{\pm.001}.$ $4.1\%$$055_{\pm.001}.$$052_{\pm.001}.$$090_{\pm.001}.$ $6.6\%$
TabSyn .$010_{\pm.000}.$$009_{\pm.000}.$$027_{\pm.001}.$ $1.5\%$$035_{\pm.001}.$$033_{\pm.002}.$$054_{\pm.001}.$ $4.1\%$
FairCB .$076_{\pm.000}.$$066_{\pm.000}.$$039_{\pm.000}.$ $6.0\%$$125_{\pm.000}.$$111_{\pm.000}.$$074_{\pm.000}.$ $10.3\%$
FairTGAN .$034_{\pm.000}.$$030_{\pm.000}.$$055_{\pm.001}.$ $4.0\%$$080_{\pm.001}.$$053_{\pm.000}.$$087_{\pm.001}.$ $7.3\%$
Ours .$126_{\pm.001}.$$121_{\pm.000}.$$109_{\pm.001}.$ $11.9\%$$201_{\pm.000}.$$174_{\pm.005}.$$174_{\pm.003}.$ $18.3\%$
Table 3: The values of machine learning efficiency with the best classifier for each dataset and DCR scores.
Adult Bank COMPAS Mean Adult Bank COMPAS Mean
[…]
Real .$928_{\pm.000}.$$936_{\pm.000}.$$810_{\pm.001}.$ $89.1\%$
CoDi .$858_{\pm.001}.$$826_{\pm.014}.$$678_{\pm.002}.$ $78.7\%$$331_{\pm.003}.$$348_{\pm.001}.$$400_{\pm.002}.$ $36.0\%$
GReaT .$901_{\pm.002}.$$688_{\pm.024}.$$717_{\pm.008}.$ $76.9\%$$320_{\pm.002}.$$348_{\pm.003}.$$377_{\pm.003}.$ $34.8\%$
SMOTE .$914_{\pm.000}.$$928_{\pm.001}.$$778_{\pm.002}.$ $87.3\%$$327_{\pm.004}.$$265_{\pm.001}.$$273_{\pm.003}.$ $28.8\%$
STaSy .$885_{\pm.003}.$$895_{\pm.003}.$$728_{\pm.013}.$ $83.6\%$$344_{\pm.001}.$$345_{\pm.002}.$$362_{\pm.001}.$ $35.0\%$
TabDDPM .$907_{\pm.001}.$$917_{\pm.002}.$$745_{\pm.001}.$ $85.6\%$$339_{\pm.000}.$$350_{\pm.002}.$$367_{\pm.005}.$ $35.2\%$
TabSyn .$911_{\pm.000}.$$919_{\pm.000}.$$749_{\pm.001}.$ $86.0\%$$339_{\pm.002}.$$351_{\pm.005}.$$367_{\pm.006}.$ $35.2\%$
FairCB .$915_{\pm.001}.$$907_{\pm.002}.$$771_{\pm.001}.$ $86.4\%$$054_{\pm.000}.$$031_{\pm.001}.$$012_{\pm.001}.$ $3.2\%$
FairTGAN .$881_{\pm.000}.$$863_{\pm.006}.$$705_{\pm.002}.$ $81.6\%$$348_{\pm.002}.$$348_{\pm.004}.$$374_{\pm.006}.$ $35.7\%$
Ours .$893_{\pm.002}.$$914_{\pm.002}.$$734_{\pm.001}.$ $84.7\%$$344_{\pm.001}.$$350_{\pm.002}.$$370_{\pm.004}.$ $35.5\%$
Table 2 and Table 3 show that our method remains competitive with SOTA models in terms of column-wise density estimation, pair-wise column correlations, and machine learning efficiency, even after incorporating a uniformly distributed direction for sensitive features during the sampling phase. Across the three experimental datasets, our model achieved an average AUC of $84.7\%$, which is only $1.3\%$ lower than TabSyn, the SOTA deep learning tabular synthesis method. Although altering the original sensitive distribution is anticipated to reduce AUC, the results demonstrate that our method effectively preserves high fidelity in the generated synthetic data. While SMOTE and FairCB excel in fidelity, their low DCR scores suggest that they closely mimic the real data, which is expected given their direct interpolation of the original data points.
5.4.2 Fairness Scores in Classification Tasks
Table 4: The values of DPR and EOR with best classifier for each dataset.
Adult Bank COMPAS Mean Adult Bank COMPAS Mean
[…]
Real .$309_{\pm.001}.$$402_{\pm.015}.$$675_{\pm.006}.$ $46.2\%$$193_{\pm.005}.$$367_{\pm.024}.$$645_{\pm.015}.$ $40.2\%$
CoDi .$293_{\pm.034}.$$189_{\pm.054}.$$855_{\pm.025}.$ $44.6\%$$247_{\pm.042}.$$172_{\pm.063}.$$857_{\pm.031}.$ $42.5\%$
GReaT .$249_{\pm.015}.$$572_{\pm.289}.$$624_{\pm.066}.$ $48.2\%$$155_{\pm.028}.$$380_{\pm.126}.$$543_{\pm.075}.$ $35.9\%$
SMOTE .$321_{\pm.006}.$$405_{\pm.028}.$$648_{\pm.021}.$ $45.8\%$$254_{\pm.011}.$$381_{\pm.026}.$$589_{\pm.009}.$ $40.8\%$
STaSy .$261_{\pm.045}.$$468_{\pm.123}.$$436_{\pm.092}.$ $38.8\%$$182_{\pm.069}.$$451_{\pm.113}.$$433_{\pm.140}.$ $35.5\%$
TabDDPM .$261_{\pm.006}.$$337_{\pm.020}.$$558_{\pm.041}.$ $38.5\%$$156_{\pm.007}.$$334_{\pm.028}.$$540_{\pm.047}.$ $34.3\%$
TabSyn .$281_{\pm.017}.$$336_{\pm.016}.$$697_{\pm.040}.$ $43.8\%$$178_{\pm.019}.$$317_{\pm.023}.$$664_{\pm.062}.$ $38.6\%$
FairCB .$286_{\pm.002}.$$719_{\pm.005}.$$675_{\pm.006}.$ $56.0\%$$192_{\pm.003}.$$801_{\pm.016}.$$638_{\pm.021}.$ $54.4\%$
FairTGAN .$554_{\pm.006}.$$338_{\pm.093}.$$448_{\pm.025}.$ $44.7\%$$697_{\pm.017}.$$158_{\pm.033}.$$392_{\pm.041}.$ $41.6\%$
Ours .$543_{\pm.026}.$$710_{\pm.057}.$$800_{\pm.080}.$ $68.4\%$$667_{\pm.059}.$$649_{\pm.040}.$$825_{\pm.113}.$ $71.4\%$
As shown in Table 4, our method significantly outperforms all baseline models, achieving a mean DPR of $68.4\%$ and a mean EOR of $71.4\%$. In comparison, state-of-the-art tabular generative models and FairTGAN fall below $50\%$ for both fairness metrics, while FairCB remains below $60\%$ for both. From Table 3, our method achieves an average AUC of $84.7$, which is higher than FairTGAN and only $1.6$ lower than FairCB. By jointly optimizing performance and fairness, our method presents a more balanced trade-off than existing approaches. Moreover, in Section E.1, we demonstrate that simply overwriting the distribution of sensitive attributes is insufficient to mitigate bias in downstream classification tasks and can lead to a decrease in classification performance.
5.4.3 Sensitive Feature Distributions
In addition to evaluating the fairness of machine learning models, we also examine the class distribution of sensitive attributes in synthetic data. To achieve this, we use stacked bar plots to visualize the contingency tables for sensitive features and the target label. We further compare the distribution of sensitive attributes between real and synthetic data using bar plots.
Figure 2: The distribution of sensitive attributes across different target label values on the Adult dataset using our method.
We present the distribution of sensitive attributes versus the target label on the Adult dataset using our method in Figure 2. The synthetic data generated by our approach demonstrates a balanced joint distribution between sensitive features and the target label, effectively reducing the disparities observed in the original data. Additionally, we provide plots of the real data distribution and synthetic distributions generated by other methods in Appendix D. Note that we only visualize the synthetic data generated by our method, FairTGAN, and FairCB, which are three fairness-aware generative models, because we assume that the other baseline models replicate the real data distribution. In comparison, our method effectively produces a more balanced joint distribution of sensitive attributes and target labels than the FairTGAN, FairCB, and the real distribution.
@@ 5.4.3 Sensitive Feature Distributions
Figure 3: Comparison in the real versus synthetic distribution of sensitive attributes across all datasets.
5.4.4 Performance-Fairness Trade-Off
Balancing fairness and predictive accuracy is a key challenge in synthetic data generation, particularly when fairness improvements might impact important outcomes. This section explores how adjustments in fairness levels influence both performance and fairness metrics in the generated data. Our method provides a way to control this trade-off, allows practitioners to find a balance that meets their specific needs, whether they prioritize fairness, accuracy, or a compromise between the two.
To analyze the trade-off between fairness and accuracy, we define a balancing level $i$ ranging from 00 to $10$. This level quantifies how much we adjust the original joint distribution of the target and sensitive attributes to make it more uniform.
Suppose we have $N$ sensitive attributes, each attribute with $C_{i}$ unique values, where $i$ ranges from 1 to $N$. This results in $K$ combinations of values where $K=C_{i}\times\cdots\times C_{N}$. Under the same targe label, each combination has a probability $y_{k}$ where $k$ is one of the $K$ combinations. For each combination, we calculate an offset $d_{k}=\bar{y}-y_{k}$, the difference between the average probability $\bar{y}={\sum_{k}y_{k}}/{|\mathbf{y}|}$ and the current probability $y_{k}$. The offset $d_{k}$ represents how much the current joint probability deviation $y_{k}$ for a combination of sensitive attributes and the target label deviate from the average probability $\bar{y}$. By identifying these deviations, we can measure how far the joint distribution is from a uniform distribution, which is often desired for achieving fairness. For a balancing level $i\in[0,10]$, the balancing ratio is $i/10$, and the adjusted values $y_{k}^{\text{balanced}}$ are calculated as:
$y_{k}^{\text{balanced}}=y_{k}+d_{k}\times\frac{i}{10}$ (8)
The balanced values $y_{k}^{\text{balanced}}$ represent the adjusted probabilities for each combination of sensitive attributes and the target label after applying a fairness transformation. The balancing process ensures that $y_{k}^{\text{balanced}}$ systematically reduces disparities in the joint distribution of sensitive attributes and target labels, moving it closer to uniformity as the balancing level $i$ increases. By choosing an appropriate balancing level, the method allows for a practical balance between performance and fairness.
The results presented in the previous sections are based on a balancing level of 10, which achieves a nearly uniform joint distribution of the target and sensitive attributes. We conducted experiments to assess how different balancing levels affect the performance-fairness trade-off. We evaluated metrics such as AUC, DPR, and EOR, with DPR and EOR averaged across all sensitive attributes. We computed a composite score as a weighted sum of these metrics, with weights of $0.5$ for AUC, $0.25$ for DPR, and $0.25$ for EOR. Figure 4 illustrates the trade-off between fairness and performance across different balancing levels. The optimal composite scores for the Adult and Bank datasets are achieved at a balancing level of $10$, indicating a uniform distribution. For the COMPAS dataset, the best score is at level $9$, representing a nearly uniform distribution. Practitioners can adjust the metric weights to suit specific application needs or establish thresholds for individual metrics based on their requirements.
Figure 4: Trade-off between performance and fairness across balancing levels for experimental datasets. SCORE was computed as a weighted sum of these metrics, with weights of 0.5 for AUC, 0.25 for DPR, and 0.25 for EOR. The best composite scores for the Adult and Bank datasets are achieved at a balancing level of 10, while the COMPAS dataset achieves the best score at level 9.
6 Limitations and Discussion
Our proposed method is designed to generate balanced synthetic tabular data while considering sensitive attributes. However, it requires sensitive features to be specified in advance, which can be challenging in large-scale enterprise datasets with thousands of features. Furthermore, we could extend our testing to datasets with more than two sensitive attributes and imbalanced target distributions.
In terms of efficiency, we use a U-Net with attention layers as the backbone neural network for the posterior estimator in the diffusion framework, but this approach is more time-consuming compared to TabSyn. As the number of sensitive attributes increases, there are more hyperparameters to tune, which makes fine-tuning more computationally demanding, especially with the long sampling time. In the future, we aim to explore methods to reduce computational costs. We believe the main bottlenecks in computational costs are due to the long diffusion timesteps and the quadratic complexity of attention modules. Possible solutions include reducing diffusion timesteps via acceleration methods such as Flow-DPM-Solver Xie et al. (2024) and refining the backbone, for instance, by employing linear attention Katharopoulos et al. (2020) or by using a pre-trained autoencoder to embed tabular data into a lower-dimensional latent space.
Lastly, our proposed method is designed to condition on specified group labels, which limits its applicability in scenarios where demographic information is unavailable or restricted due to privacy concerns. Future iterations of the model could integrate privacy-preserving techniques, such as synthetic attribute imputation, to enable fairness-aware data generation without requiring explicit sensitive attribute disclosure.
On the other hand, our proposed method generates fair data with a balanced joint distribution of the target label and multiple sensitive attributes, achieving higher fairness scores than baseline methods. Additionally, the model shows strong potential for extending to multimodal data synthesis. If we can improve the efficiency of our model and successfully adapt it to multimodal scenarios, we believe it could be widely applied to address fairness issues in fields such as healthcare, finance, and beyond.